OpenStudy (anonymous):

what is the slope of -48=2x-8y

OpenStudy (anonymous):

the answer has to be in fraction form

OpenStudy (anonymous):

Turn the equation into function form: -48=2x-8y 8y=2x+48 /:8 y=1/4x+6 slope: 1/4

OpenStudy (nincompoop):

does it? in the slope y-intercept form \[y = mx + b\] m is the slope